OpenStudy (anonymous):

A winery has a cylindrical fermentation tank 20 feet high with a radius of 6 feet. If the tank is being filled with grape juice at a rate of 90 cubic feet per minute, how fast is the surface of the juice rising at any time t ?

OpenStudy (anonymous):

\[V=\pi r^2h\] You're given that \(\dfrac{dV}{dt}=90\) and you must find \(\dfrac{dh}{dt}\). Differentiating with respect to \(t\), you have \[\frac{dV}{dt}=2\pi r\frac{dr}{dt}h+\pi r^2\frac{dh}{dt}\] Notice that as the level of the juice rises, the radius does not change, so \(\dfrac{dr}{dt}=0\): \[\frac{dV}{dt}=\pi r^2\frac{dh}{dt}\] Plug in the known values: \[90=\pi 6^2\frac{dh}{dt}~~\Rightarrow~~\frac{dh}{dt}=\frac{90}{36\pi}\approx0.796\text{ ft/min}\]